OpenStudy (anonymous):

show that each pair is a solution of the equation 2y+x=5;(-1,3) (7,-1)

OpenStudy (anonymous):

just put values of x & y in your equation and verify that your statement is still true

OpenStudy (anonymous):

i dnt understand were to put them

OpenStudy (anonymous):

In an ordered pair (-1,3) the first number is the x and the second is the y (x,y) so you put -1 in for x and 3 in for y 2(3) + (-1) = 5 6 + (-1) = 5 5 = 5 so that point (-1,3) is on the line 2x + y = 5 now try the next point and I will check you out.

OpenStudy (anonymous):

2(7)+-1=5 did i write that right

OpenStudy (anonymous):

OK.. the point is (7,-1) so x = 7 and y = -1 your equation is 2y + x = 5 make sure you put in x = 7 and y = -1

OpenStudy (anonymous):

so this isnt a coordinate

OpenStudy (anonymous):

2(-1) + 7 = 5 -2 + 7 = 5 5 = 5 so (7,-1) is also on that line... it is a solution.

OpenStudy (anonymous):

yes... it is a coordinate.. but you put the x value of 7 into the y variable and you put the y value of -1 into the x variable. that is the mistake you made.

OpenStudy (anonymous):

i put them in the wrong place

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

i got what ur saying so it would always be in the order when i fill in xand y